OpenStudy (ray10):

Use a characteristic equation to find the particular solution of \[\large \frac{ d^{2} x}{ dt^{2} }-4\frac{ dx }{ dt }-32x=0\] given that x(0)=5 and x'(0)=16

OpenStudy (ray10):

so far worked out that the homogeneous comes to\[\huge Ae^{-4x}+Be^{8x}\] Can someone please check that and help me with the rest?

terenzreignz (terenzreignz):

You seem familiar, have we encountered each other on here before?

OpenStudy (ray10):

hahahha sarcasm? :P

terenzreignz (terenzreignz):

Decide for yourself. Right now however, we are faced with a rather daunting differential equations problem. Well, YOU'RE faced with it, I may well be only here to watch you... >:)

OpenStudy (ray10):

well I'm hoping you remember me :O yes rather daunting :/ that is true :(

terenzreignz (terenzreignz):

Well, I remember people based on feats... you managed a fiasco with Pan involving complex roots... am I right?

OpenStudy (ray10):

ah yes you are very right!

terenzreignz (terenzreignz):

Now this appears to be nothing more than an initial value problem, are you really daunted by this? :D

OpenStudy (ray10):

I have a large disliking for differential equations requiring this :(

terenzreignz (terenzreignz):

I see. What was your name again?

OpenStudy (ray10):

Ray

terenzreignz (terenzreignz):

Oh.. right, silly me :) TJ. Terry. Terence. Pleasure's mine. Now, let's get to work... I want to make sure you knew what you were doing here, would you mind telling me how you arrived at that general solution you posted immediately after your question?

OpenStudy (ray10):

I remember you saying to call you TJ.! :) Glad to be talking! well first I went to find the homogeneous solution by solving the first eqution;

OpenStudy (ray10):

equation*

terenzreignz (terenzreignz):

Go on...

OpenStudy (ray10):

\[\large \lambda^{2}-4\lambda-32=0\] then it can be found that \[\large \lambda =-4\] or \[\large \lambda =8\]

OpenStudy (ray10):

then I looked at the table and the example I have is similar, using the homogeneous solution as \[\large Ae^{\alpha*x}+Be^{\beta*x}\]

terenzreignz (terenzreignz):

I see. Well I can tell you that the general solution you posted is incorrect, however :)

terenzreignz (terenzreignz):

Much as I'd like to see you flail about in search of an error, even I'm not THAT sadistic. The error was minor, but an error nonetheless: \[\huge x(t)=Ae^{-4\color{red}t}+Be^{8\color{red}t}\]

OpenStudy (ray10):

I see :/ could you please walk me through where I went wrong? I'm not to bright with using the table for this :(

terenzreignz (terenzreignz):

Relax. Minor error. Refer to previous post ^

OpenStudy (ray10):

Oh!!! Right! Of course! thank you :)

OpenStudy (ray10):

I didn't see that myself

terenzreignz (terenzreignz):

Now, to the "hard" part, the initial values. First, the initial values not only involve x(t) but also x'(t) Care to find x'(t) ?

OpenStudy (ray10):

this is where I seem to get lost, in previous questions, the right hand side had another expression

OpenStudy (ray10):

this one has a zero

terenzreignz (terenzreignz):

Again, relax. It's not nearly as hard as you think. \[\huge x(t)=Ae^{-4t}+Be^{8t}\]\[\huge x'(t) = \frac{d}{dt}x(t)\] in other words, just differentiate the bloody solution XD

OpenStudy (ray10):

ohhh differentiate the homogeneous equation! :O

terenzreignz (terenzreignz):

Oh, the sound of realisation... I get that often ;)

OpenStudy (ray10):

I am very well aware you do! :P \[\large 8*Be^{8t}-4*Ae^{-4t}\]

terenzreignz (terenzreignz):

Now, let me recap for you, ok? \[\Large x(t) = Ae^{-4t}+Be^{8t}\]\[\Large x'(t) = 8Be^{8t}-4Ae^{-4t}\] catch me so far?

OpenStudy (ray10):

yes I sure do

terenzreignz (terenzreignz):

Good. This entire thing is just a fancy way to present a system of linear equations. I trust you'd have no problem with that? .... I didn't think you would. NOW. x(0) = 5 How do you put that to use?

OpenStudy (ray10):

oh I see, so now it's just the simple use of: \[\large 5=Ae^{-4*0}+Be^{8*0}\]

OpenStudy (ray10):

\[\large 5=Ae^{0}+Be^{0}\]

terenzreignz (terenzreignz):

And even simpler, A + B = 5 Good. Now put x'(0) = 16 to use.

OpenStudy (ray10):

\[\large -4A+8B=16\] also \[\large -4\]

OpenStudy (ray10):

don't worry about the second negative four

terenzreignz (terenzreignz):

A+B = 5 -4A + 8B = 16 Hardly daunting anymore, aye?

OpenStudy (ray10):

not at all anymore! solve simultaneously :D

terenzreignz (terenzreignz):

I shall wait.

OpenStudy (ray10):

work out A and B and then substitute them into the equation! :D \[\large 2e^{-4t}+3e^{8t}\]

terenzreignz (terenzreignz):

That wasn't too bad, now was it?

OpenStudy (ray10):

It made so much sense when you helped me see that it's a simple question just differently worded, thank you !!! :D

terenzreignz (terenzreignz):

I am... TERENZ LOL Signing off for now ^_^ --------------------------- Terence out

OpenStudy (ray10):

Ter-iffic!!